# A Generalized Control Function Approach to Production Function Estimation

### v1

@@ 2 Results
holds at the true production function. Moreover, if a production function satisfies moment condition (5), then it satisfies moment condition (4) for the control function $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$. Finally, if $\varphi(z_{it})$ and $q_{it}-f(k_{it},v_{it})$ have finite $L^{2}$ norms, then moment condition (5) is Neyman orthogonal with respect to $L^{2}$-integrable perturbations of $\left(\tilde{\varphi}_{it},\tilde{q}_{it}-\tilde{f}_{it}\right)$.
@@ 4 Concluding Remarks
Johns Hopkins University November 26, 2025
@@
Next, we prove that moment condition (5) implies moment condition (4) with $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$. Note that
@@
where the first equality is due to the law of iterated expectations and the second equality to $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}[q_{it}-f(k_{it},v_{it})|z_{it}\setminus{z^{\dagger}_{it}},\text{E}[q_{it-1}|x_{it-1}]]=0$ by the definition of $\tilde{q}_{it}-\tilde{f}_{it}$. Adding moment condition (5) and moment condition (Proof of Theorem 2.) implies moment condition (4) with $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$.

### v2

@@ 2 Results
holds at the true production function. Moreover, if a production function satisfies moment condition (5), then it satisfies moment condition (4) for the control function $h(z_{it}\setminus z^{\dagger}_{it}$, $\text{E}[q_{it-1}|x_{it-1}])=\tilde{q}_{it}-\tilde{f}_{it}$. Finally, if $\varphi(z_{it})$ and $q_{it}-f(k_{it},v_{it})$ have finite $L^{2}$ norms, then moment condition (5) is Neyman orthogonal with respect to $L^{2}$-integrable perturbations of $\left(\tilde{\varphi}_{it},\tilde{q}_{it}-\tilde{f}_{it}\right)$.
@@ 4 Concluding Remarks
Johns Hopkins University November 27, 2025
@@
Next, we prove that moment condition (5) implies moment condition (4) with $h(z_{it}\setminus z^{\dagger}_{it}$, $\text{E}[q_{it-1}|x_{it-1}])=\tilde{q}_{it}-\tilde{f}_{it}$. Note that
@@
where the first equality is due to the law of iterated expectations and the second equality to $\tilde{q}_{it}-\tilde{f}_{it}=\text{E}[q_{it}-f(k_{it},v_{it})|z_{it}\setminus{z^{\dagger}_{it}},\text{E}[q_{it-1}|x_{it-1}]]=0$ by the definition of $\tilde{q}_{it}-\tilde{f}_{it}$. Adding moment condition (5) and moment condition (4) implies moment condition (4) with $h\left(z_{it}\setminus z^{\dagger}_{it},\text{E}[q_{it-1}|x_{it-1}]\right)=\tilde{q}_{it}-\tilde{f}_{it}$.